Number 4
Draw a net for the solid. Then find the surface area. Round to the nearest whole number.It is shaped like a triangle which that is shaped like a tent. The number beside it are 12ft, 14ft, 26ft, 14ft. With 5 sides total.
1
Expert's answer
2012-06-07T08:00:25-0400
S=((a+b)/2)*sqrt(c^2-(((b-a)^2+c^2-d^2)/2(b-a))^2)=((12+14)/2*sqrt(26^2-(((14-12)^2+26^2-14^2)/2(14-12))^2)=13*sqrt(676-((4+676-196)/4)^2)=13*sqrt(442335)=8646